OpenStudy (anonymous):

Solve 3x2 = 12x. @hartnn

hartnn (hartnn):

ok, first subtract 12x from both sides, what u get ?

OpenStudy (anonymous):

-9x^2 = x

hartnn (hartnn):

it doesn't work that way, 3x^2 and 12x are not like terms, so u cannot combine them subtracting 12x from both sides will give you \(3x^2-12x=0\) got this ?

OpenStudy (anonymous):

Oops >_< Okay

hartnn (hartnn):

then factor out 'x' from \(3x^2-12x\) can you ?

OpenStudy (anonymous):

Basically factor it all, or just the x's?

OpenStudy (uri):

3x2 -12x=0 3x(x-4)

hartnn (hartnn):

you can factor 3x from 3x^2-12x like for example, factoring 5x from 15x^2-25x will give you 5x (3x-5)

hartnn (hartnn):

you understand factoring ?

OpenStudy (anonymous):

Yes, I'll do that right now.

OpenStudy (anonymous):

3x ( x-4)

OpenStudy (uri):

Yess :p

hartnn (hartnn):

yeah, so 3x(x-4) = 0 gives 3x=0 x-4 = 0 what 2 values of x you get from these ?

OpenStudy (anonymous):

Don't worry, I saw your answer, but I still worked it out as well @uri @hartnn -4 and 3?

hartnn (hartnn):

3x= 0 gives you x= 0, right! x-4 = 0 adding 4 on both sides, x-4+4 = 0+4 x= +4 got these ?

OpenStudy (anonymous):

Yes, thank you!

hartnn (hartnn):

welcome ^_^